# Novel Tour Construction Heuristic for Pick-Up and Delivery Routing Problems

### v1

@@ Abstract
Index Terms:
Traveling salesman problems, Vehicle routing and navigation, Intelligent Transportation Systems, Algorithms, Autonomous robot, Autonomous vehicles
I Introduction
@@
Given a set of locations, the Precedence Constrained Traveling Salesperson Problem (TSP-PC) finds the shortest possible tour that respects precedence relations between locations, visiting each location exactly once and returning to the starting location [1]. These problems arise in a variety of real-world applications such as crane scheduling in ports, job scheduling, automotive paint shops, helicopter scheduling and flexible manufacturing systems [2, 3, 4, 5], where commodities have to be efficiently transported between locations to delivery locations, and a tour is only feasible when pick up locations are visited before their respective deliveries.
@@
Exact methods to solve TSP-PCs are limited to problems with a relatively small number of locations, typically below 100, due to their $\mathcal{NP}$-hard nature [6, 7, 8]. Conversely, heuristics follow tour construction rules to iteratively select locations, defining their order of visitation. While tour construction heuristics generate feasible solutions quickly, they do not guarantee optimality and are generally sensitive to problem parameters. For this reason, heuristics are typically used when solutions have to be found instantaneously [9, 10, 11], or to initialize exact methods for faster convergence to the optimal solution [12].
@@
Due to the presence of precedence constraints, sophisticated heuristics tailored for the TSP-PC have been neglected in literature [13, 14], and instead, the simple Nearest Neighbor (NN) greedy heuristic is commonly used [15, 16]. When precedence constraints do not exist, the Convex Hull Cheapest Insertion (CHCI) heuristic has been shown to produce superior solutions when compared to the NN heuristic [17, 18]. The CHCI heuristic is initiated by a subtour created from the convex hull of points, and its interior points are then progressively added to the subtour in increasing order of insertion cost ratios until the complete tour is obtained. However, the CHCI heuristic has not been adapted to the TSP-PC because of the challenges in accounting for precedence constraints when inserting points to a subtour.
@@
The main contribution of this paper is the extension of the CHCI algorithm to the TSP-PC by initiating a subtour using the convex hull boundary points of only the non-delivery locations. After reordering the subtour to originate at the depot, the subtour is then explored in both clockwise and counterclockwise directions because precedence constraints introduce a sense of direction of travel. For every ensuing insertion, the adapted CHCI (ACHCI) heuristic maintains precedence constraints by only permitting insertions into feasible subtour segments. This implies that pickup locations can be added anywhere in the subtour, while delivery locations can only be added in the subtour segment that has already visited its corresponding pickup location. Test cases that replicate real-world pickup and delivery considerations are created by adding precedence constraints to the TSPLIB benchmark instances [19]. After comparing the performance of the ACHCI and NN heuristics, their solutions are used to warm start an exact solver to study convergence.
II Precedence Constrained TSP
@@
$\sum_{j\in\mathcal{V}}x_{ij}y_{jm}=y_{im}+\sum_{j\in\mathcal{V}}x_{ij}q_{jm}~{}\forall m\in\mathcal{H},j\in\mathcal{V}$ (1i)
@@
The problem of optimally visiting package pickup and delivery locations is shown in Fig. 1, where, depending on the application, $C_{ij}$ may capture distance, operational, fuel or any expense to be minimized between two locations $i,j\in\mathcal{V}$. For an intracity problem, if $C_{ij}$ denotes distance, the tour cost of a truck that starts at the depot, picks up and delivers all the packages and returns to the depot typically amounts to tens of miles. When applied to helicopter scheduling, this may rise to hundreds of miles [2], while in the context of flexible manufacturing systems, it is often measured in feet [4].
To generalize the analysis of the proposed ACHCI heuristic for any TSP-PC application, the popular TSPLIB benchmark instances [19] are used with added precedence constraints, and the total Euclidean distance is minimized in this paper. TSPLIB instance ‘eil51’ is shown in Fig. 2, where precedence constraints are illustrated using grey line segments that relate parent pickup positions, shown in red, with the respective children drop-off positions, marked in green. The agent starts and ends its tour at the depot, marked in blue.
III Adapted CHCI Heuristic
@@
If the identically located depot node $\{0,2n+1\}\notin T_{0}$, consecutive nodes $v_{i}^{*},v_{j}^{*}$ with the lowest insertion cost ratio is first found in $T_{0}$, and the depot node is inserted between them. To start and end the subtour at the depot, the updated subtour $T_{0}$ is reordered to $[0,v_{j}^{*},...,v_{i}^{*},2n+1]$. Because a sense of direction exists due to the precedence constraints, denote the resulting subtour as $\overleftarrow{T}$. The subsequent steps will be repeated in the other direction $\overrightarrow{T}$ as well, and the lower cost tour is selected.
@@
IV Nearest Neighbor Heuristic
The benchmark for the ACHCI algorithm is the NN which is known to produce feasible tours in real time for problems with a large number of locations [20]. The NN heuristic uses a fast greedy selection rule and is commonly seen in constrained TSP literature because it can easily be modified to account for constraints in the problem formulation [21, 15]. Starting from the depot, the NN heuristic assigns the nearest unvisited node as the next node until all nodes are contained in the tour. By modifying this to only consider unvisited feasible nodes as the next node, it can be adapted for the TSP-PC:
@@ IV Nearest Neighbor Heuristic
V Computational Experiments
To compare the effectiveness of the ACHCI heuristic with the NN heuristic, sufficiently diverse benchmark instances are not readily available for the TSP-PC. For this reason, the popular TSPLIB benchmark instances [19] are modified in a reproducible manner to create cases precedence constraints.
@@ V Computational Experiments
V-A Comparison of Heuristic Solutions
The ACHCI and NN tours are compared for 60 TSPLIB instances with precedence constraints, with the number of points varying from 51 to 1,577. The result of these experiments is summarized in Table II and Table II for ten instances. The first column lists the name of each TSPLIB instance, formatted with a prefix followed by a numeric value that indicates the number of points in the respective instance. The second and third columns provide the solution costs of the NN and ACHCI heuristics along with the percentage reduction in NN cost by using the ACHCI solution. The histogram in Fig. 4c shows that the ACHCI tour cost is lower than the NN cost in all cases with centrally located children. However, the ACHCI heuritic does not perform as well when parent nodes are closer to the centroid, as seen in Fig. 4d.
TABLE I: Centrally Located Children
@@ V-A Comparison of Heuristic Solutions
TABLE II: Centrally Located Parents
@@
V-B Heuristic Initialization of an Exact Algorithm
@@
The MILP reformulation of Eq. (1) is modeled in a Matlab R2023b environment and the intlinprog solver of Gurobi 9.5.2 [22] is utilized on an AMD Ryzen 5600X CPU clocked at 3.7 GHz, paired with 128GB RAM. The time limit of the exact solver is set to 5 minutes and the results are tabulated in Tables II and II. The fourth column provides the solution found when running the exact algorithm standalone without any heuristic initialization, where it is seen that for instances with 99 or more points, no feasible solution is found within the allocated computation time. The fifth and sixth columns show the solution found when the exact solver uses the NN solution and ACHCI solution respectively for a warm start, along with the percentage reduction in solution cost compared with the standalone solution.
V-C Discussion
@@
When the ACHCI heuristic is used to initialize an exact solver, it consistently improves convergence of the exact solver in all of the cases studied. The initiation of the candidate subtour with the convex hull of the points is advantageous when precedence constraints do not exist, because points on the boundary of the convex hull are visited in the same cyclic order as their order in the optimal tour [23]. The ACHCI heuristic makes use of this property and therefore, the resulting solution acts as a good initialization point for the exact solver. In most NN warm start cases as well, the convergence of the exact solver was improved.
VI Conclusion
In this paper, the popular CHCI algorithm has been adapted for the TSP-PC, a route optimization problem commonly seen in logistic applications where items have to be picked up from and delivered to various locations. The proposed ACHCI algorithm is a tour construction heuristic that accounts for precedence constraints by first drawing the initiating convex hull over non-delivery locations and permitting subsequent insertions only in feasible segments that respect precedence constraints.
Given the diverse problem parameters and cost considerations in real-world TSP-PC instances, the TSPLIB benchmark dataset was used to generalize the performance analysis of the ACHCI heuristic. Using the ACHCI heuristic solution to initialize an exact solver improved convergence in all cases, while the NN heuristic solutions did not always improve convergence.
The ACHCI heuristic solution outperforms the NN heuristic in all cases where delivery locations are centrally located, though it did not perform as well otherwise. It is therefore highly applicable to spatial configurations of logistics operations where pickup locations are situated around the periphery to facilitate efficient collection and delivery locations are centrally situated for sorting, processing, or distribution. Examples include layouts seen in flexible manufacturing systems or warehousing facilities where inventory shelves are at the perimeters of the facility, or in online retail or digital commerce settings where sorting facilities are located on the outskirts of a city and customers are within city premises.

### v3

@@ Abstract
keywords:
Traveling salesman problems, Algorithms, Logistics, Heurisitcs
[…]
latexText page 5 contains only floats
1 Introduction
@@
Given a set of locations, the Precedence Constrained Traveling Salesperson Problem (TSP-PC) finds the shortest possible tour that respects precedence relations between locations, visiting each location exactly once and returning to the starting location (Mingozzi et al., 1997). These problems arise in a variety of real-world applications such as tool path optimization, switching energy minimization of program compilers, automotive paint shops, and flexible manufacturing systems (Kucukoglu et al., 2019; Shobaki and Jamal, 2015; Spieckermann et al., 2004), where commodities have to be efficiently transported between pickup locations to delivery locations, and a tour is only feasible when pick up locations are visited before their respective deliveries.
@@
Exact methods to solve TSP-PCs are limited to problems with a relatively small number of locations, typically below 100, due to their $\mathcal{NP}$-hard nature (Jamal et al., 2017; Shobaki and Jamal, 2015; Salii, 2019). Conversely, heuristics follow tour construction rules to greedily select locations, defining their order of visitation. While tour construction heuristics generate feasible solutions quickly, they do not guarantee optimality and are generally sensitive to problem parameters. For this reason, heuristics are typically used when solutions have to be found instantaneously (Xiang et al., 2008; Wong et al., 2014; Marković et al., 2015), or to initialize exact methods for faster convergence to the optimal solution (Braekers et al., 2014).
@@
Due to the presence of precedence constraints, sophisticated heuristics tailored for the TSP-PC have been neglected in literature (Glover et al., 2001; Taillard, 2022), and instead, the simple Nearest Neighbor (NN) greedy heuristic is commonly used (Bai et al., 2020; Edelkamp et al., 2018). When precedence constraints do not exist, the Convex Hull Cheapest Insertion (CHCI) heuristic has been shown to produce superior solutions when compared to the NN heuristic (Ivanova et al., 2021; Warburton, 1993). The CHCI heuristic is initiated by a subtour created from the convex hull of points, and its interior points are then progressively added to the subtour in increasing order of insertion cost ratios until the complete tour is obtained. However, the CHCI heuristic has not been adapted to the TSP-PC because of the challenges in accounting for precedence constraints when inserting points to a subtour.
@@
The main contribution of this paper is the extension of the CHCI algorithm to the TSP-PC by initiating a subtour using the convex hull boundary points of only the non-delivery locations. After reordering the subtour to originate at the depot, the subtour is then explored in both clockwise and counterclockwise directions because precedence constraints introduce a sense of direction of travel. For every ensuing insertion, the adapted CHCI (ACHCI) heuristic maintains precedence constraints by only permitting insertions into feasible subtour segments. This implies that pickup locations can be added anywhere in the subtour, while delivery locations can only be added in the subtour segment that has already visited its corresponding pickup location. Test cases that replicate real-world pickup and delivery considerations are created by adding precedence constraints to the TSPLIB benchmark instances (Reinhelt, 2014). After comparing the performance of the ACHCI and NN heuristics, their solutions are used to warm start an exact solver to study convergence.
2 Precedence Constrained TSP
@@
$\sum_{j\in\mathcal{V}}x_{ij}y_{jm}=y_{im}+\sum_{j\in\mathcal{V}}x_{ij}q_{jm}~{}\forall m\in\mathcal{H},i\in\mathcal{V}$ (1i)
@@
The problem of optimally visiting package pickup and delivery locations is shown in Fig. 1, where, depending on the application, $C_{ij}$ may capture distance, operational, fuel or any expense to be minimized between two locations $i,j\in\mathcal{V}$. For an intracity problem, if $C_{ij}$ denotes distance, the tour cost of a truck that starts at the depot, picks up and delivers all the packages and returns to the depot typically amounts to tens of miles. When applied to helicopter scheduling, this may rise to hundreds of miles (Fiala Timlin and Pulleyblank, 1992), while in the context of flexible manufacturing systems, it is often measured in feet (Ascheuer et al., 1993).
To generalize the analysis of the proposed ACHCI heuristic for any TSP-PC application, the popular TSPLIB benchmark instances (Reinhelt, 2014) are used with added precedence constraints, and the total Euclidean distance is minimized in this paper. TSPLIB instance ‘eil51’ is shown in Fig. 2, where precedence constraints are illustrated using grey line segments that relate parent pickup positions, shown in red, with the respective children drop-off positions, marked in green. The agent starts and ends its tour at the depot, marked in blue.
3 Adapted CHCI Heuristic
@@
If the identically located depot node $\{0,2n+1\}\notin T_{0}$, consecutive nodes $v_{i}^{*},v_{j}^{*}$ with the lowest insertion cost ratio are first found in $T_{0}$, and the depot node is inserted between them. To start and end the subtour at the depot, the updated subtour $T_{0}$ is reordered to $[0,v_{j}^{*},...,v_{i}^{*},2n+1]$. Because a sense of direction exists due to the precedence constraints, denote the resulting subtour as $\overleftarrow{T}$. The subsequent steps will be repeated in the other direction $\overrightarrow{T}$ as well, and the lower cost tour is selected.
@@
4 Nearest Neighbor Heuristic
The benchmark for the ACHCI algorithm is the NN which is known to produce feasible tours in real time for problems with a large number of locations (Charikar et al., 1997). The NN heuristic uses a fast greedy selection rule and is commonly seen in constrained TSP literature because it can easily be modified to account for constraints in the problem formulation (Grigoryev and Tashlykov, 2018; Bai et al., 2020). Starting from the depot, the NN heuristic assigns the nearest unvisited node as the next node until all nodes are contained in the tour. By modifying this to only consider unvisited feasible nodes as the next node, it can be adapted for the TSP-PC:
@@ 4 Nearest Neighbor Heuristic
5 Computational Experiments
To compare the effectiveness of the ACHCI heuristic with the NN heuristic, sufficiently diverse benchmark instances are not readily available for the TSP-PC. For this reason, the popular TSPLIB benchmark instances (Reinhelt, 2014) are modified in a reproducible manner to create cases with precedence constraints.
@@ 5 Computational Experiments
5.1 Comparison of Heuristic Solutions
The ACHCI and NN tours are compared for 60 TSPLIB instances with precedence constraints, with the number of points varying from 51 to 1,577. The result of these experiments is summarized in Table 1 where the first column lists the name of each TSPLIB instance, formatted with a prefix followed by a numeric value that indicates the number of points in the respective instance. The second and third columns provide the solution costs of the NN and ACHCI heuristics along with the percentage reduction in NN cost by using the ACHCI solution. The histogram in Fig. 4c shows that the ACHCI tour cost is lower than the NN cost in all cases with centrally located children. However, the ACHCI heuritic does not perform as well when parent nodes are closer to the centroid, as seen in Fig. 4d.
Table 1: TSP-PC Tour Costs
TSPLIB Centrally Located Children Centrally Located Parents
instance NN ACHCI NN ACHCI
eil51 5.8e+02 4.8e+02 (-17.7%) 5.8e+02 6.2e+02 (+ 6.3%)
t70 9.1e+02 7.9e+02 (-13.3%) 1.1e+03 1.2e+03 (+10.6%)
eil76 7.9e+02 6.3e+02 (-20.4%) 7.3e+02 8.3e+02 (+13.3%)
erlin52 1.1e+04 8.3e+03 (-26.1%) 1.1e+04 1.2e+04 (+ 5.5%)
eil101 9.7e+02 7.5e+02 (-23.0%) 9.0e+02 9.1e+02 (+ 0.7%)
rat99 1.8e+03 1.5e+03 (-17.4%) 1.8e+03 1.9e+03 (+ 7.7%)
pr76 2.0e+05 1.3e+05 (-35.0%) 1.5e+05 1.7e+05 (+11.1%)
roC100 3.5e+04 2.5e+04 (-30.4%) 2.9e+04 3.7e+04 (+27.1%)
roD100 3.5e+04 2.8e+04 (-21.2%) 2.8e+04 3.5e+04 (+26.6%)
roE100 3.3e+04 2.8e+04 (-17.2%) 3.2e+04 3.6e+04 (+12.3%)
roA100 3.0e+04 2.7e+04 (-10.2%) 3.4e+04 3.3e+04 (- 3.5%)
roB100 3.7e+04 2.7e+04 (-26.1%) 3.3e+04 3.8e+04 (+13.8%)
in105 2.6e+04 1.7e+04 (-36.7%) 2.3e+04 2.8e+04 (+20.9%)
pr107 7.5e+04 7.2e+04 (- 4.9%) 7.2e+04 8.6e+04 (+19.0%)
pr124 1.0e+05 8.1e+04 (-22.1%) 9.5e+04 1.1e+05 (+11.3%)
roB150 4.4e+04 3.3e+04 (-25.8%) 3.9e+04 4.0e+04 (+ 1.7%)
roA150 4.1e+04 3.3e+04 (-21.2%) 4.0e+04 3.9e+04 (- 1.5%)
pr136 1.5e+05 1.2e+05 (-19.8%) 1.5e+05 1.4e+05 (-10.7%)
pr144 9.4e+04 7.2e+04 (-23.3%) 7.6e+04 1.0e+05 (+35.8%)
pr152 1.1e+05 8.6e+04 (-18.5%) 1.0e+05 1.4e+05 (+43.1%)
rat195 3.2e+03 2.9e+03 (-10.4%) 3.3e+03 3.4e+03 (+ 3.4%)
bier127 1.8e+05 1.4e+05 (-25.7%) 1.7e+05 1.7e+05 (+ 4.2%)
roA200 5.6e+04 3.8e+04 (-32.3%) 4.9e+04 5.0e+04 (+ 4.1%)
roB200 4.9e+04 3.6e+04 (-27.3%) 4.1e+04 5.2e+04 (+27.8%)
rd100 1.5e+04 1.0e+04 (-32.8%) 1.2e+04 1.1e+04 (- 5.5%)
gil262 3.8e+03 3.1e+03 (-19.7%) 3.5e+03 4.2e+03 (+21.1%)
pr226 1.2e+05 1.1e+05 (- 5.8%) 1.1e+05 1.4e+05 (+21.4%)
a280 4.0e+03 3.3e+03 (-17.8%) 4.0e+03 3.7e+03 (- 5.8%)
ts225 2.1e+05 1.5e+05 (-27.8%) 1.7e+05 2.4e+05 (+46.5%)
pr264 7.4e+04 7.4e+04 (- 0.5%) 6.8e+04 7.6e+04 (+11.2%)
tsp225 6.0e+03 4.3e+03 (-27.9%) 5.2e+03 6.5e+03 (+26.0%)
pr299 8.1e+04 6.1e+04 (-24.1%) 7.0e+04 6.9e+04 (- 0.9%)
in318 6.2e+04 5.2e+04 (-16.3%) 6.8e+04 8.2e+04 (+20.6%)
in318 6.2e+04 5.2e+04 (-16.3%) 6.8e+04 8.2e+04 (+20.6%)
h130 9.3e+03 7.7e+03 (-16.9%) 9.5e+03 1.0e+04 (+ 8.0%)
u159 6.5e+04 5.4e+04 (-16.8%) 6.0e+04 8.4e+04 (+39.2%)
h150 9.5e+03 8.0e+03 (-15.3%) 9.3e+03 1.3e+04 (+34.4%)
d198 2.6e+04 1.8e+04 (-30.5%) 2.3e+04 2.1e+04 (-11.1%)
pr439 1.8e+05 1.3e+05 (-27.0%) 1.5e+05 1.5e+05 (- 3.4%)
rat575 1.1e+04 8.2e+03 (-24.1%) 1.0e+04 1.1e+04 (+ 8.9%)
rat783 1.5e+04 1.0e+04 (-29.2%) 1.3e+04 1.5e+04 (+15.8%)
rd400 2.4e+04 1.8e+04 (-23.9%) 2.1e+04 2.7e+04 (+28.9%)
fl417 1.9e+04 1.7e+04 (-11.0%) 1.8e+04 2.3e+04 (+29.4%)
pcb442 7.7e+04 5.9e+04 (-22.7%) 8.7e+04 8.2e+04 (- 5.0%)
d493 5.0e+04 4.1e+04 (-18.0%) 4.7e+04 5.0e+04 (+ 5.1%)
pr1002 4.1e+05 3.1e+05 (-24.3%) 4.0e+05 4.4e+05 (+10.1%)
u574 5.6e+04 4.5e+04 (-20.6%) 5.9e+04 6.9e+04 (+16.9%)
p654 6.8e+04 5.3e+04 (-22.5%) 5.4e+04 6.7e+04 (+23.7%)
d657 8.1e+04 5.7e+04 (-28.8%) 7.0e+04 8.5e+04 (+20.6%)
u724 6.9e+04 5.2e+04 (-25.0%) 5.8e+04 7.2e+04 (+24.7%)
u1060 3.6e+05 2.7e+05 (-25.9%) 3.4e+05 3.9e+05 (+15.2%)
vm1084 4.2e+05 2.8e+05 (-31.5%) 3.8e+05 4.0e+05 (+ 3.6%)
nrw1379 8.0e+04 6.6e+04 (-17.0%) 8.2e+04 8.9e+04 (+ 9.4%)
pcb1173 9.7e+04 7.1e+04 (-26.9%) 8.7e+04 8.5e+04 (- 2.0%)
d1291 7.7e+04 6.5e+04 (-16.1%) 8.8e+04 9.3e+04 (+ 6.3%)
rl1304 4.7e+05 3.1e+05 (-33.9%) 4.7e+05 5.4e+05 (+13.9%)
rl1323 5.8e+05 3.5e+05 (-40.1%) 4.6e+05 5.6e+05 (+21.2%)
fl1400 4.3e+04 2.9e+04 (-31.7%) 3.1e+04 3.5e+04 (+12.6%)
u1432 2.3e+05 1.8e+05 (-20.1%) 2.2e+05 2.4e+05 (+ 9.0%)
fl1577 3.8e+04 2.6e+04 (-30.1%) 3.4e+04 4.8e+04 (+38.6%)
Table 2: TSP-PC Tour Costs for Instances with Centrally Located Children
@@ 5.1 Comparison of Heuristic Solutions
Table 3: TSP-PC Tour Costs for Instances with Centrally Located Parents
@@
5.2 Heuristic Initialization of an Exact Algorithm
@@
The MILP reformulation of Eq. (1) is modeled in a Matlab R2023b environment and the intlinprog solver of Gurobi 9.5.2 (Gurobi Optimization, LLC, 2023) is utilized on an AMD Ryzen 5600X CPU clocked at 3.7 GHz, paired with 128GB RAM. The time limit of the exact solver is set to 5 minutes and the results are tabulated in Tables 3 and 3. The fourth column provides the solution found when running the exact algorithm standalone without any heuristic initialization, where it is seen that for instances with 99 or more points, no feasible solution is found within the allocated computation time. The fifth and sixth columns show the solution found when the exact solver uses the NN solution and ACHCI solution respectively for a warm start, along with the percentage reduction in solution cost compared with the standalone solution.
5.3 Discussion
@@
The initiation of the candidate subtour with the convex hull of the points is advantageous when precedence constraints do not exist, because points on the boundary of the convex hull are visited in the same cyclic order as their order in the optimal tour (Deineko et al., 1994). The ACHCI heuristic makes use of this property and therefore, the resulting solution acts as a good initialization point for the exact solver. This is seen in Tables 3 and 3, where the convergence of the exact solver is improved in every instance where the ACHCI tour is used as an initialization. In most NN warm start cases as well, the convergence of the exact solver was improved.
6 Conclusion
The proposed ACHCI algorithm is a tour construction heuristic that accounts for precedence constraints by first drawing the initiating convex hull over non-delivery locations and permitting subsequent insertions only in feasible segments that respect precedence constraints. Using the ACHCI heuristic solution to initialize an exact solver improved convergence in all cases, while the NN heuristic solutions did not always improve convergence. The ACHCI heuristic solution outperforms the NN heuristic in all cases where delivery locations are centrally located, though it did not perform as well otherwise. It is therefore highly applicable to spatial configurations of operations where pick-up locations are situated around the periphery to facilitate efficient collection and delivery locations are centrally situated for sorting, processing, or distribution. Examples include layouts in flexible manufacturing systems or warehousing facilities where inventory shelves are at the perimeters of the facility, or in online retail and digital commerce settings where sorting facilities are located on the outskirts of a city and customers are within city premises.